OpenStudy (toneva):

what is the solution set for 7x^(-2)+26x^(-1)=8

OpenStudy (3mar):

May I help?

OpenStudy (welshfella):

Multiply each term by x^2

OpenStudy (3mar):

Is that your equation? \[\large 7x^{-2}+26x^{-1}=8\] \[\large \frac{ 7 }{ x^2} +\frac{ 26 }{x} =8\]

OpenStudy (toneva):

@3mar Yes, can you please help me solve this?

OpenStudy (3mar):

That is with my pleasure! What about multiplying both sides by \(x^2\)?

OpenStudy (toneva):

@3mar Would cross out the X under the 26 too? 7+26x=8x^2 or would it still be 7+26x^2/x=8x^2

OpenStudy (3mar):

\[\Huge \color{CadetBlue }{7+26x=8x^2 }\] That is very good! You are Awesome!

OpenStudy (3mar):

Can you put this equation in the standard form?